1.     A firm manufactures two products; the net profit on product 1 is Rupees 3 per unit and Rupees 5 per unit on product 2. The manufacturing process is such that each product has to be processed in two departments D1 and D2. Each unit of product1 requires processing for 1 minute at D1 and 3 minutes at D2; each unit of product 2 requires processing for 2 minutes at D1 and 2 minutes at D2. Machine time available per day is 860 minutes at D1 and 1200 minutes at D2. How much of product 1 and 2 should be produced every day so that total profit is maximum. Make the mathematical model for the given problem.



1
Expert's answer
2021-10-13T17:32:37-0400

Let "x_1" and "x_2" be levels of production of two products, then

"z(x_1,x_2)=3\cdot x_1+5\cdot x_2"  is profit function and it should be maximized so "3\cdot x_1+5\cdot x_2\rarr max"

For manufacturing  "x_1"unis of product 1 and "x_2" units of product 2 we need in

using "x_1+2\cdot x_2" minutes on depatment D1 and "3\cdot x_1+2\cdot x_2" minutes on department D2 that have availiable machine time 800 and 1200 minutes per day correspondingly therefore we have two restrictions for the plan

"x_1+2\cdot x_2\le800;\\
3\cdot x_1+2\cdot x_2\le1200;\\"

We also must take in account positiveness restrictions :

"x_1\ge0,x_2\ge0"

Thus we have math model of production:

"3\cdot x_1+5\cdot x_2\rarr max"

"x_1+2\cdot x_2\le800;\\
3\cdot x_1+2\cdot x_2\le1200;\\"

"x_1\ge0,x_2\ge0"

Standard form convenient for simplex method is

z->max;

"z-3\cdot x_1-5\cdot x_2=0;"

"x_1+2\cdot x_2+1\cdot t+0\cdot s=800;\\
3\cdot x_1+2\cdot x_2+0\cdot t+1\cdot s=1200;\\
x_1,x_2,t,s\ge0"

To solve we use simplex method:

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 z & x_1 & x_2 &t&s&b\\ \hline
 1 & -3 & -5 &0&0&0\\
 \hdashline
 0 & 1 & ! 2&1&0&800\\
 \hdashline
 0 & 3 & 2&0&1&1200\\
\end{array}"


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 z & x_1 & x_2 &t&s&b\\ \hline
 1 & -1/2 & 0 &5/2&0&2000\\
 \hdashline
 0 & 1/2 & 1&1/2&0&400\\
 \hdashline
 0 & ! 2 & 0&-1&1&400\\
\end{array}"


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 z & x_1 & x_2 &t&s&b\\ \hline
 1 & 0 & 0 &9/4&1&2100\\
 \hdashline
 0 & 0 & 1&3/4&-1&300\\
 \hdashline
 0 & 1 & 0&-1/2&1/2&200\\
\end{array}"


Answer:

"x_1=200,x_2=300,z_{max}=2100"